# Image-Text-Image Knowledge Transferring for Lifelong Person Re-Identification with Hybrid Clothing States

### v1

@@ Abstract.
With the continuous expansion of intelligent surveillance networks, lifelong person re-identification (LReID) has received widespread attention, pursuing the need of self-evolution across different domains. However, existing LReID studies accumulate knowledge with the assumption that people would not change their clothes. In this paper, we propose a more practical task, namely lifelong person re-identification with hybrid clothing states (LReID-Hybrid), which takes a series of cloth-changing and cloth-consistent domains into account during lifelong learning. To tackle the challenges of knowledge granularity mismatch and knowledge presentation mismatch that occurred in LReID-Hybrid, we take advantage of the consistency and generalization of the text space, and propose a novel framework, dubbed $Teata$, to effectively align, transfer and accumulate knowledge in an “image-text-image” closed loop. Concretely, to achieve effective knowledge transfer, we design a Structured Semantic Prompt (SSP) learning to decompose the text prompt into several structured pairs to distill knowledge from the image space with a unified granularity of text description. Then, we introduce a Knowledge Adaptation and Projection strategy (KAP), which tunes text knowledge via a slow-paced learner to adapt to different tasks without catastrophic forgetting. Extensive experiments demonstrate the superiority of our proposed $Teata$ for LReID-Hybrid as well as on conventional LReID benchmarks over advanced methods.
Person Re-Identification, Lifelong Learning, Hybrid Clothing States
ccs: Computing methodologies Object identification
Figure 1. Conventional lifelong ReID task (LReID) assumes pedestrians would always exhibit the same clothing in the lifelong evolution process of ReID models. To promote ReID applications, we propose a more practical task by taking the clothing changes into account. It is much more challenging since the mismatch of knowledge caused by the hybrid clothing states across domains can lead to learning confusion and exacerbate knowledge forgetting. “SC” and “CC” denote same-cloth and cloth-changing states, respectively.
1. Introduction
@@
Lifelong Person Re-identification (LReID) is a sub-field of person ReID, which aims to retrieve the same individual that appeared in different camera views, while incrementally accumulating the ReID knowledge in different scenarios (Pu et al., 2021). This topic offers numerous meaningful application scenarios, capable of accommodating the ongoing expansions of surveillance networks and facilitating continual learning of ReID models across diverse domains. Therefore, it has increasingly captivated attention and research efforts. For example, Ge et al. (Ge et al., 2022) replay old samples saved in memory with specific-designed supervisions to balance stability and plasticity for lifelong learning. Yu et al. (Yu et al., 2023) further introduce knowledge refreshing and consolidation techniques to achieve positive forward and backward transfer. To avoid data privacy, Lu et al. (Lu et al., 2022) present dream memory to preserve knowledge, while learning incrementally with augmented geometric distillation.
Despite advanced progress in these efforts, existing LReID approaches unfortunately ignore another factor that is common and of concern in the ReID application scenario, i.e., clothing changes (Gu et al., 2022; Qian et al., 2020; Jin et al., 2022; Yang et al., 2019). For example, models could be easily deceived by people with clothing changes (CC) even after sufficient steps of continual learning on conventional data, since the acquired knowledge has the bias that the same identity keeps the same clothes (SC). To overcome this limitation, we propose a more practical LReID task, namely lifelong person re-identification with hybrid clothing states (LReID-Hybrid), which requires constantly updating models on a series of cloth-changing and conventional ReID datasets. As depicted in Fig. 1, the goal of LReID-Hybrid is to improve the model’s ability to generalize across various scenarios with different clothing conditions.
@@
However, developing a lifelong ReID model on sequential datasets with clothing state differences is a non-trivial task and inevitably encounters several challenges. The first one is knowledge representation mismatch. Although the incremental datasets all belong to the task of person ReID, they naturally have distribution differences due to locations and times, which causes visual representation shifts between different learning steps. This mismatch is much worse by changes in the state of people’s clothing. Clothing features are viewed as a discriminative representation in conventional ReID datasets, but this representation is disruptive when the same person changes their clothes (i.e., cloth-changing scenarios). Such a representation mismatch further exacerbates knowledge catastrophic forgetting, leading to fluctuations in the performance of the model during lifelong learning. The second challenge is knowledge granularity mismatch. In contrast to conventional ReID, cloth-changing ReID typically requires models to perceive more fine-grained cues in person images (Wang et al., 2023b), since visual appearance information, which takes up a greater proportion, can be conversely misleading. The difference in representation granularity is data-driven and has no impact on learning in a single data domain. Nevertheless, it occurs when carrying lifelong learning with multiple datasets, which inevitably hinders effective knowledge transfer and accumulation.
@@
In light of the great success of visual language models in effective representation (Radford et al., 2021), as well as the advanced applications in ReID (Li et al., 2023), we seek a more compact and robust representation of person images in the text space. Natural language can intuitively and easily describe various features of person images, including appearance, gender, behavior, etc, some of which are conducive to the re-identification of person identity. More importantly, natural language can use a unified sentence pattern or vocabulary when describing person images, which can effectively bypass the domain differences existing in image pixels. Based on the above inspiration, we present a novel framework, dubbed Teata, that takes full advantage of Text modality to align, transfer and accumulate knowledge. Our core idea is a two-stage process of alternating iterations, to first distill pixel-level information and category labels into text semantic embeddings, and then project image features onto the text space to guide the image encoder to acquire ReID knowledge. Similar to CLIP-ReID (Li et al., 2023), we utilize CLIP technology to align image and text features, thereby transferring the accumulated knowledge in the text space.
@@
To complement the text modality, we further propose the Structured Semantic Prompt (SSP) learning and the Knowledge Adaptation and Projection (KAP) strategy to tackle the aforementioned challenges. Concretely, we use learnable text prompts as a container for image-to-text distillation in the first stage, and decompose the prompt into several structured pairs. Each pair contains a set of tokens shared across all tasks, designed to learn representation objects of person images, and a specific set of tokens for learning identity-related image representation content. With the explicit structured design, we can transfer knowledge between different tasks with the uniform granularity of text descriptions. To alleviate the knowledge representation mismatch, we additionally tune the distilled text embeddings of different identities in the text semantic space during the second-stage learning. We optimize it with a slow-paced learner, so as to avoid knowledge confusion or forgetting and adjust for differences in representation between different tasks. Benefiting from the generalization of the text representation and the proposed novel modules, our Teata is rehearsal-free and capable of preserving old knowledge while adapting to new tasks for lifelong ReID of hybrid clothing conditions.
Contributions. We summarize the key contributions as follows,
@@
(1) We exploit a more practical lifelong ReID task, namely LReID with hybrid clothing states (LReID-Hybrid). It requires incrementally learning a ReID model under different clothing states. To this end, we present a novel framework Teata to realize knowledge alignment, transferring and accumulation in an “image-text-image” closed loop.
@@
(2) To alleviate the limitation of knowledge granularity mismatch, we propose the structured semantic prompt learning to explicitly decompose text prompts into a series of structured tokens. It unifies the granularity of text representation, thus facilitating the knowledge transfer.
@@
(3) We introduce a knowledge adaptation and projection strategy to optimize text representations of every identity via a slow-paced learner. It can not only adapt to the representation differences of different tasks when guiding learning of image encoders, but also avoid knowledge forgetting concisely and effectively.
@@
(4) Extensive experiments demonstrate the superiority of our proposed Teata for LReID-Hybrid over various baselines. Additionally, it is capable of achieving state-of-the-art results with significant margins on conventional LReID benchmarks.
2. Related Work
2.1. Person Re-Identification
@@
There has been remarkable progress in person re-identification leveraging pre-prepared stationary training data (Wang et al., 2018; Luo et al., 2019; Zheng et al., 2019; Wang et al., 2023a). Conventionally, researchers focus on the same-cloth scenario where people would always show a relatively stable clothing appearance. Based on this assumption, to reduce the influence of intra-class variations, Zheng et al. (Zheng et al., 2019) propose to leverage the generated data and Zhou et al. (Zhou et al., 2019) propose to perform omni-scale feature learning. Wang et al. (Wang et al., 2018) design a multi-branch deep network architecture integrating discriminative information with various granularities. Some methods turn to auxiliary data clues (Jin et al., 2020; He et al., 2021) for help.
@@
Although researchers have made efforts to promote the discriminative ability of ReID models, conventional person ReID models would fail when pedestrians change their clothes, since they almost exclusively focus on clothes. Xu et al. (Xu et al., 2021) and Eom et al. (Eom et al., 2021) attempt to use generative models to augment samples by explicitly synthesizing person images with various clothes, so as to learn more robust features against clothing changes. Some studies draw support from auxiliary modalities to assist learning soft-biometrics features unrelated to clothes (Qian et al., 2020; Yang et al., 2019; Wang et al., 2022; Jin et al., 2022; Hong et al., 2021; Chen et al., 2021; Liu et al., 2023; Guo et al., 2023; Agbodike et al., 2023; Cui et al., 2023). For instance, SPT+ASE (Yang et al., 2019) utilizes reliable and discriminative curve patterns on the body contour sketch. CAMC (Wang et al., 2022) uses heatmaps of human postures to encode body shape semantic information. GI-ReID (Jin et al., 2022) learns cloth-agnostic representations by leveraging personal unique and cloth-independent gait information. Other researchers take advantage of clothing labels to eliminate the negative effects of clothing features. For instance, CAL (Gu et al., 2022) adopts adversarial learning to penalize the model’s predictive power to clothes. AIM (Yang et al., 2023) adopts a dual-branch model to simulate causal intervention and eliminate clothing bias. However, simply ignoring the clothes of pedestrians would hurt the performance in the same-cloth scenario, in which clothing can be one of the most discriminative characteristics.
2.2. Lifelong Person Re-Identification
@@
Lifelong learning seeks to maintain stable performance on old tasks while adapting the models to gain new knowledge (Li and Hoiem, 2017; Wu et al., 2019). Due to the expansion of smart surveillance systems, person ReID models also have to continually accumulate knowledge information from old domains and generalize well on new domains (Pu et al., 2021). To meet the demand, AKA (Pu et al., 2021) adopts knowledge distillation baselines (Li and Hoiem, 2017) to preserve acquired knowledge and maintains a learnable knowledge graph to adaptively update previous knowledge. However, due to significant domain variations, it struggles to retain old knowledge without access to previous data. GwFReID (Wu and Gong, 2021) formulates a comprehensive learning objective for maintaining coherence during progressive learning. Recently, rehearsal-based methods have achieved state-of-the-art results by saving a few exemplars in each domain. For instance, PTKP (Ge et al., 2022) proposes a pseudo task knowledge preservation framework to alleviate the domain gap in the last BN layer. KRC (Yu et al., 2023) introduces a dynamic memory model for bi-directional knowledge transfers and a knowledge consolidation scheme. However, due to privacy issues, it is impractical to save person images. As is opposite to reality, all previous works assume each pedestrian would always show a cloth-consistent appearance in the lifelong evolution process. They overlook the influence of clothing state variation and tend to fail when clothing information is not useful to distinguish different people. Differently, we explore a more practical lifelong setting where both cloth-consistent domains and cloth-changing domains are considered. We show that the consistency, compactness, and generalization of text semantics can help integrate different knowledge across domains with various clothing states.
@@
Figure 2. Framework of our proposed Teata. It starts by aligning image features with text embeddings to transfer visual information and category labels to the text space. Next, we leverage the robustness and consistency of text semantics to guide the discriminative feature learning of visual representations. To achieve effective knowledge transfer, we design a Structured Semantic Prompt (SSP) learning in the first stage, which decomposes the original text prompt into several structured pairs to distill knowledge from the image space with a unified granularity of text description. Then, we introduce a Knowledge Adaptation and Projection strategy (KAP) in the second stage, which tunes text knowledge via a slow-paced learner to adapt to different tasks without catastrophic forgetting.
3. Methodology
3.1. Formulation and Overview

### v2

# Image-Text-Image Knowledge Transfer for Lifelong Person Re-Identification with Hybrid Clothing States

@@ Abstract
With the continuous expansion of intelligent surveillance networks, lifelong person re-identification (LReID) has received widespread attention, pursuing the need of self-evolution across different domains. However, existing LReID studies accumulate knowledge with the assumption that people would not change their clothes. In this paper, we propose a more practical task, namely lifelong person re-identification with hybrid clothing states (LReID-Hybrid), which takes a series of cloth-changing and same-cloth domains into account during lifelong learning. To tackle the challenges of knowledge granularity mismatch and knowledge presentation mismatch in LReID-Hybrid, we take advantage of the consistency and generalization capabilities of the text space, and propose a novel framework, dubbed Teata, to effectively align, transfer, and accumulate knowledge in an “image-text-image” closed loop. Concretely, to achieve effective knowledge transfer, we design a Structured Semantic Prompt (SSP) learning to decompose the text prompt into several structured pairs to distill knowledge from the image space with a unified granularity of text description. Then, we introduce a Knowledge Adaptation and Projection (KAP) strategy, which tunes text knowledge via a slow-paced learner to adapt to different tasks without catastrophic forgetting. Extensive experiments demonstrate the superiority of our proposed Teata for LReID-Hybrid as well as on conventional LReID benchmarks over advanced methods.
Figure 1: Comparison between the conventional LReID task and our proposed LReID with hybrid clothing states (LReID-Hybrid). Person images from the same dataset with the same bounding box colour share the same identity. “SC” and “CC” denote domains with same-cloth and cloth-changing states, respectively. Emojis represent the attitudes of the model towards acquiring new knowledge. Conventional lifelong ReID task (LReID) assumes there are no clothing changes for each person throughout the lifelong evolution process. This allows shared appearance knowledge to be acquired for ReID across different domains, and thus the process of acquiring new knowledge is joyful. We propose a more practical task that takes the clothing changes into account. The mismatch of knowledge caused by varying clothing states across domains can exacerbate knowledge forgetting, making the process of acquiring new knowledge confusing.
I Introduction
@@
Lifelong Person Re-identification (LReID) is a sub-field of person ReID, which aims to retrieve the same individual that appeared in different camera views, while incrementally accumulating the ReID knowledge in different scenarios [1]. This topic offers numerous meaningful application scenarios, capable of accommodating the ongoing expansion of surveillance networks and facilitating the continual learning of ReID models across diverse domains. Consequently, it has increasingly captivated attention and research efforts. For example, Ge et al. [2] replay old samples saved in memory with specific-designed supervisions to balance stability and plasticity for lifelong learning. Yu et al. [3] further introduce knowledge refreshing and consolidation techniques to achieve positive forward and backward transfer. To avoid data privacy, Lu et al. [4] present dream memory to preserve knowledge, while learning incrementally with augmented geometric distillation.
Under the efforts of existing work, shared appearance knowledge for ReID can be accumulated across different domains in the conventional LReID setting, as shown in Fig. 1 (a). Unfortunately, they overlook a crucial factor that is common and of concern in the ReID application scenario, i.e., clothing changes [5, 6, 7, 8]. They follow the same-cloth ReID setting, where each person always wears the same clothes, so the acquired knowledge is mostly related to clothing appearances. When pedestrians change their clothes, the ReID models will struggle to achieve satisfactory results [9, 5, 10]. Similarly, in the lifelong learning scenario, the assumption about the clothing state is not conducive to the generalization and robustness of the model. For example, models could be easily deceived by people with clothing changes (CC) even after sufficient steps of continual learning on conventional same-cloth data, since the acquired knowledge has the bias that the same identity keeps the same clothes (SC).
Therefore, we advocate taking the clothing changes into account in LReID. Considering that other coexisting factors, such as camera view changes and background changes, are still thorny and cannot be ignored, we formulate a challenging LReID setting using datasets with different clothing states based on existing LReID research [1, 2, 3]. Specifically, we propose a more practical LReID task, namely lifelong person re-identification with hybrid clothing states (LReID-Hybrid), which requires constantly updating the model on a series of cloth-changing and conventional same-cloth ReID datasets, as depicted in Fig. 1 (b). In the LReID-Hybrid setting, the model is expected to resolve the conflict of knowledge acquired when the clothing states of data alternate, and accumulate shareable and complementary knowledge from hybrid data that comes separately and incrementally. Eventually, the learned model is required to be generalized across various scenarios with different clothing conditions.
@@
However, developing a lifelong ReID model on sequential domains with clothing state differences is a non-trivial task and inevitably encounters several challenges. The first one is knowledge representation mismatch. Although the incremental datasets all belong to the task of person ReID, they naturally exhibit distribution differences due to variations in location and time, causing visual representation shifts between different learning stages. This mismatch is exacerbated by changes in the state of people’s clothing. Clothing features, which serve as discriminative representations in conventional ReID datasets, become disruptive when the same person changes clothes (i.e., cloth-changing scenarios). Such a representation mismatch further exacerbates catastrophic forgetting of knowledge, leading to fluctuations in the performance of the model during lifelong learning. The second challenge is knowledge granularity mismatch. In contrast to conventional ReID, cloth-changing ReID typically requires models to perceive more fine-grained cues in person images [10], since visual appearance information, which takes up a greater proportion, can be conversely misleading. The difference in representation granularity is data-driven and has no impact on learning in a single data domain. Nevertheless, it occurs when carrying lifelong learning across hybrid domains, which inevitably hinders effective knowledge transfer and accumulation. The above two challenges can be especially tricky when data with hybrid clothing states comes separately and sequentially in the lifelong learning setting.
@@
In light of the great success of visual language models in effective representation [11], as well as their advanced applications in ReID [12], we thus seek a more compact and robust representation of person images in the text space. Natural language can intuitively and easily describe various features of person images, including appearance, gender, behavior, and more, some of which are conducive to re-identifying individuals. More importantly, natural language employs a unified sentence pattern or vocabulary when describing person images, effectively bypassing the domain differences inherent in image pixels. Based on the above inspiration, we present a novel framework, dubbed Teata, that takes full advantage of Text modality to align, transfer and accumulate knowledge. Our core idea is a two-stage process of alternating iterations, to first distill pixel-level information and category labels into text semantic embeddings, and then project image features onto the text space to guide the image encoder in acquiring ReID knowledge. Similar to CLIP-ReID [12], we utilize CLIP technology to align image and text features, thereby transferring the accumulated knowledge into the text space.
@@
To complement the text modality, we further propose the Structured Semantic Prompt (SSP) learning and the Knowledge Adaptation and Projection (KAP) strategy to tackle the aforementioned challenges of knowledge mismatches in granularity and representation, respectively. Concretely, we use learnable text prompts as containers for image-to-text distillation in the first stage, and decompose the prompt into several structured pairs. Each pair contains a set of tokens shared across all tasks for learning representation objects of person images, and a specific set of tokens for learning identity-related image representation content. With the explicit structured design, we can transfer knowledge between different tasks with uniform text description granularity. To alleviate the knowledge representation mismatch, we additionally tune the distilled text embeddings of different identities in the text semantic space during the second-stage learning. We optimize it with a slow-paced learner, so as to avoid knowledge confusion or forgetting and adjust for differences in representation between different tasks. Benefiting from the generalization of text representation and the proposed novel modules, our Teata is rehearsal-free and capable of preserving old knowledge while adapting to new tasks for lifelong ReID of hybrid clothing conditions.
Contributions. We summarize our contributions as follows,
@@
(1) We exploit a more practical lifelong ReID task, namely LReID with hybrid clothing states (LReID-Hybrid). It requires incrementally learning a ReID model under varying clothing states. To this end, we present a novel framework Teata, which facilitates knowledge alignment, transfer, and accumulation in an “image-text-image” closed loop.
@@
(2) To address the limitation of knowledge granularity mismatch, we propose the Structured Semantic Prompt learning. It explicitly decomposes text prompts into a series of structured tokens, unifying the granularity of text representation and thereby enhancing knowledge transfer.
@@
(3) We introduce a Knowledge Adaptation and Projection strategy to optimize the text representations of each identity using a slow-paced learner. This strategy not only adapts to the representation differences across tasks when guiding the learning of image encoders but also prevents knowledge forgetting concisely and effectively.
@@
(4) Extensive experiments demonstrate the superiority of our proposed Teata framework for LReID-Hybrid over various baselines. Additionally, it achieves state-of-the-art results with significant margins on conventional LReID benchmarks.
II Related Work
II-A Person Re-Identification
@@
There has been remarkable progress in person re-identification leveraging pre-prepared stationary training data [13, 14, 15, 16]. Conventionally, researchers focus on the same-cloth scenario where people would always show a relatively stable clothing appearance. Based on this assumption, to reduce the influence of intra-class variations, Zheng et al. [15] propose to leverage the generated data and Zhou et al. [17] propose to perform omni-scale feature learning. Wang et al. [13] design a multi-branch deep network architecture integrating discriminative information with various granularities. Some methods turn to auxiliary data clues [18, 19] for help. Wang et al. [16] rethink the role of the classifier as an interactive ReID process and propose a strong baseline from a projection-on-prototypes perspective. Considering the huge annotation cost of identity labels, unsupervised person ReID has also received widespread attention [20, 21, 22]. Following a clustering-based training strategy, Liu et al. further propose a stochastic learning strategy and propose a unified distance matrix during clustering, and Lin et al. generate style-transferred training images with different camera styles to achieve across-camera similarity exploration.
@@
Although researchers have made efforts to promote the discriminative ability of person ReID models [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], these models would still fail when pedestrians change their clothes, since they almost exclusively focus on clothes. Xu et al. [23] and Eom et al. [24] attempt to use generative models to augment samples by explicitly synthesizing person images with various clothes, so as to learn more robust features against clothing changes. Some studies draw support from auxiliary modalities to assist learning soft-biometrics features unrelated to clothes [6, 8, 9, 7, 25, 26, 27, 28, 29, 30]. For instance, SPT+ASE [8] utilizes reliable and discriminative curve patterns on the body contour sketch. CAMC [9] uses heatmaps of human postures to encode body shape semantic information. GI-ReID [7] learns cloth-agnostic representations by leveraging personal unique and cloth-independent gait information. Other researchers take advantage of clothing labels to eliminate the negative effects of clothing features. For instance, CAL [5] adopts adversarial learning to penalize the model’s predictive power to clothes. AIM [31] adopts a dual-branch model to simulate causal intervention and eliminate clothing bias. Considering using well-trained off-the-shelf tools for modality extraction or manually annotated clothing labels can be inflexible, some studies [10, 32] focus on learning robust representations for cloth-changing person ReID by leveraging the mined fine-grained attributes [10] and abundant semantics [32] from images.
Recent research has successfully developed models trained on either cloth-changing or same-cloth datasets. These models are capable of recognizing different pedestrians in scenarios with a consistent clothing state. However, they struggle to perform well across domains with different clothing states, as knowledge acquired from cloth-changing and same-cloth domains exhibits significant differences in representation and granularity. For instance, ignoring cloth-related information would harm performance in same-cloth scenarios, where clothing can be one of the most discriminative characteristics. The relatively coarse-grained knowledge acquired from the same-cloth domain struggles to help distinguish pedestrians in cloth-changing scenarios.
II-B Lifelong Person Re-Identification
@@
Lifelong learning seeks to maintain stable performance on old tasks while adapting the models to gain new knowledge [33, 34]. Due to the expansion of smart surveillance systems, person ReID models also have to continually accumulate knowledge information from old domains and generalize well on new domains [1]. To meet the demand, AKA [1] adopts knowledge distillation baselines [33] to preserve acquired knowledge and maintains a learnable knowledge graph to adaptively update previous knowledge. However, due to significant domain variations, it struggles to retain old knowledge without access to previous data. GwFReID [35] formulates a comprehensive learning objective for maintaining coherence during progressive learning. To further promote LReID, ranking consistency distillation [36], distribution-aware prototypes [37], and long short-term knowledge consolidation [38] are proposed. Recently, rehearsal-based methods have achieved state-of-the-art results by saving a few exemplars in each domain. For instance, PTKP [2] proposes a pseudo task knowledge preservation framework to alleviate the domain gap in the last Batch Normalization layer. KRC [3] introduces a dynamic memory model for bi-directional knowledge transfers and a knowledge consolidation scheme. However, it is impractical to save person images due to privacy issues, and the distribution discrepancy in the image space can aggravate knowledge forgetting [39].
Additionally, as is opposite to reality, all previous works assume each pedestrian would always show a cloth-consistent appearance in the lifelong evolution process. They overlook the influence of clothing state variation and tend to fail when clothing information is not useful to distinguish different people. Differently, we explore a more practical lifelong setting where both same-cloth domains and cloth-changing domains are considered. We show that the consistency, compactness, and generalization of text semantics can help integrate different knowledge across domains with various clothing states. Our proposed method does not rely on replayed exemplars and relieves the drawback of accumulating knowledge only in the image space.
@@
Figure 2: Framework of our proposed Teata. With the $t$-th training dataset $\mathcal{D}^{(t)}_{train}$, we first transfer visual information and category labels to the text space by aligning the image representations extracted by the image encoder $\mathcal{G}_{t-1}$ and the text embeddings extracted by the text encoder $\mathcal{T}$. To achieve effective knowledge transfer, we design the Structured Semantic Prompt learning. It uses structured prompts to distill knowledge from the image space with a unified granularity of text description. Next, the distilled text embeddings are used to guide the evolution of the image encoder $\mathcal{G}_{t}$. For effective evolution, we introduce the Knowledge Adaptation and Projection strategy to tune text knowledge via a slow-paced learner and adapt $\mathcal{G}_{t}$ without catastrophic forgetting.
III Methodology
III-A Formulation and Overview
III-A1 Task Definition